cerna.gif (1167 bytes) List of Actions from MPP Meetings lhc_logo.gif (4175 bytes)

MPP

Description

Responsible

Comments

169th (14.09.2018)
  • Test steps of the emergency dump procedure related to the LBDS after TS2.

M. Solfaroli & N. Magnin

169th (14.09.2018)
  • Request functionality to open up excitation window of ADT and prepare OP sequence for scraping the full LHC beam.

D. Valuch, MPP

163rd (27.04.2018)
  • prepare a summary slide to show which parameters are checked and how, referring to the AOB - Abort Gap Keeper - protection of fine delay parameter.

Jorg

161st (23.03.2018)
  • the TDI pressure interlock should be masked with the new expert mode and unmasked when needed; the injection veto interlock will not be touched.

Jorg

161st (23.03.2018)
  • verify with MP3 if a SIS interlock on DFB temperature is required or if it is sufficient to take care of this via procedures.

MPP

160th (16.03.2018)
  • send schematics of interlock to MPE to verify the interlock card and if re-testing is required.

A. Rossi

158th (23.02.2018)
  • check on redundant powering of the LBDS (UPS test).

ABT

After the meeting, Chiara gave an update, stating that the test was pending until CV had completed some interventions.
157th (26.01.2018)
  • Request from the impedance team an official statement on the use of the crystal chamber with ion physics beam.

Collimation

 
157th (26.01.2018)
  • Include into the BLM disabling rules a restriction for IPQs in LSS that BLMs in position 2 and 3 are not disabled at the same time.

E.B. Holzer

 
156th (08.12.2017)
  • Define use cases for the DOROS BPMs in IR7 collimators.

CollWG

Ongoing
155th (24.11.2017)
  • Operational scenarios and bunch intensities, beam types to be considered for the new BPMs electronics for IR6.

OP, MPP

Ongoing
152nd (22.09.2017)
  • Verify the phase advances from the MKDs to the TCPs and how many bunches can escape the TCDQ if injected the abort gap.

TE-ABT, Collimation

An MD (2930) was performed, results are still to be presented back at the MPP.
151st (01.09.2017)
  • Summarize planned changes in the BLMs around the TDEs in an ECR.

TE-ABT, BE-BI

 
151st (01.09.2017)
  • Verify expected beat beating at critical elements like TCTs or IR6 dump protection in case of a CLIQ mis-firing.

Bjorn

Ongoing
Joint MPP & CWG (24.07.2017)
  • Investigate whether it is needed to implement a smoother change of the tolerance during squeeze instead of going from 4 to 1 sigmas at the end. Roderik commented that the beam size is not linear with beta* so going with sqrt(beta*) might be more relevant.

OP, Coll

Will be presented by Jorg in MPP158.
149th (23.06.2017)
  • Review other failure cases for RD1 and estimate the associated FMCM signals.

MPE-PE

Simulations of the signals due to an inter-turn short in a D1 magnet are ongoing.
146th (12.05.2017)
  • Study the feasibility of a variable window for PC interlocking.

EPC

 
145th (21.04.2017)
  • Produce procedure to be performed after any change of the AGK length.

N. Magnin

 
143rd (31.03.2017)
  • BLM inhibit in case of too high transverse losses with injection: propose commissioning test to verify functionality.

BE-BI, TE-ABT

 
143rd (31.03.2017)
  • BCCM: Make post mortem data operational for BCCM and reduce BCCM trigger levels for all running sums during the beam commissioning to fully exploit the experience from loss maps etc.

BE-BI

Pending, the equipment changed ownership.
139th (03.02.2017)
  • Provide input to Christos concerning testing the BLM reaction time at 450 GeV and top energy in case of a RD1, RD34 powering failure.

MPP

The answer will be provided next week.
137th (09.12.2016)
  • Organise a circulating checklist with points to be followed up between the CC team and MPP; iterate during a coffee discussion early January.

CC team, MPP

The ECR document is still to be submitted.
133rd (16.09.2016)
  • Verify which vacuum valves closed during the 2/9/16 MKI erratic.

TE-ABT

 
133rd (16.09.2016)
  • Further understand the quench of Q2 during the 2/9/16 MKI erratic.

BLMTWG, MP3

Is the focus of a specific working group. An MD (#2930) was performed to reproduce the conditions leading to the quench.
125th (22.04.2016)
  • Discuss how to check the status of the BCT buffer. Evaluate the new IQC module to see how often the SPS BCT buffers jump to zero.

Mirko

To be presented in MPP during the EYETS.
124th (08.04.2016)
  • Confirm number of extra bunches the IP7 collimators can take during a type 2 erratic and if this causes a problem.

R. Bruce

Pending.
119th (23.10.2015)
  • Discuss the implementation of a passive beam presence flag
  • Conduct systematic simulation studies on beam scraping by collimators in case of non-working dump

M. Solfaroli, M. Valette

An EoF Md on beam scraping in case of non-working dump was conducted in 2017, an operational procedure is being prepared.
118th (09.10.2015)
  • Include cross-check of BIS self-triggering in BIS IPOC module

I.Romera

Wil be done in 2017.
104th (13.03.2015);
  • Discuss the final thresholds for collimators including BLM response between BLMTWG and the Collimation team

B. Auchmann, S. Redaelli

Final threshold update to take place in ealry 2016
101th (30.01.2015);
  • Finalize a list of commissioning steps for the BCCM

D. Belohrad

Ongoing.
98th (31.10.2014)
  • Communicate flag to be surveyed by the LHC SIS for the remote surveillance of the VME and cPCI crates to Jorg and agree on testing.

N. Magnin, E.Carlier, J.Wenninger

Still not available.
93rd (27.06.2014)
  • Calculate magnets damage limits at 450GeV and 7TeV.

A. Lechner, MPP

PhD thesis of Vivien Raginel finished. PhD of Andreas Will ongoing.
92nd (23.05.2014)
  • Review protection limit of inner triplet (80K) with realistic impact distributions.

COLL, MPP

Pending.

 


Last updated 11-Oct-2018 17:40